OpenStudy (anonymous):

Mr. Robertson was looking at the test grades of two of his students, shown in the table below. Test 1 Test 2 Test 3 Test 4 Jimmy 83 91 90 88 Brian 78 82 70 70 Which statement below best describes the students' grades? Jimmy's median test score is less than Brian's median test score by 12 points. Jimmy's mean test score is greater than Brian's mean test score by 13 points. Jimmy's mean test score is less than Brian's mean test score by 13 points. Jimmy's median test score is greater than Brian's median test score by 14 points.

OpenStudy (texaschic101):

Jimmy -- 83,88,90,91 mean = (83 + 88 + 90 + 91)/4 = 352/4 = 88 median (middle number) = (88 + 90) / 2 = 178/2 = 89 Brian -- 70,70,78,82 mean = (70 + 70 + 78 + 82)/4 = 300/4 = 75 median = (70 + 78)/2 = 148/2 = 74 can you find the answer from here ?

OpenStudy (texaschic101):

check out the mean scores

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

Its B

OpenStudy (texaschic101):

yep...you got it :)

OpenStudy (anonymous):

Thanks for takeing the time to explain

OpenStudy (anonymous):

:) :)

OpenStudy (texaschic101):

no problem...thats what I am here for :)